### v3

@@ 8.1 Pushing Gaia beyond its limits to assist in asteroseismology of intermediate- and high-mass stars
Pushing even more beyond limits, we can now fully integrate the knowledge from Gaia’s 3-band and Kepler’s (or in the future TESS’s and PLATO’s) single-band light curves to improve asteroseismic modelling. Some of the future capacities of this synergistic integration is illustrated in Fig. 11, which compares the amplitudes of the dominant dipole prograde mode in the Gaia BP and Kepler bands for 309 Dor stars. This result is extremely encouraging, because theory predicts the amplitudes to be higher at the BP than at the wavelength for such a mode in Dor pulsators (Aerts et al. 2004). While Gaia’s time-series photometry may suffer from instrumental frequencies at (sub-)multiples of the satellite frequency near 4 d, the signal in the BP band is not hampered by background noise for amplitudes above 4 mmag (Aerts et al. 2023). Predictions for amplitude ratios derived from the theory of non-radial oscillations (see Sect. 6.1 in Aerts et al. 2010) are in agreement with the results shown in the bottom panel in Fig. 11. This illustrates that Gaia also has a role to play in asteroseismic modelling, while it was not at all made for such science case. Indeed, the bottom panel shows the amplitude ratios of the 23 Dor stars among the sample of 309, for which a harmonic fit with the dominant Kepler frequency explains at least 50% of the overall variance in the Gaia BP and RP time series. Measured amplitudes depend on the inclination angle of the star’s symmetry axis of the oscillations with respect to the line-of-sight and on the position of the surface nodal lines. However, both dependencies drop out of the theoretical expression of amplitude ratios with respect to a reference wavelength for slow rotators (see Chapter 6, Aerts et al. 2010, for details). Theoretical expressions for the ratios in the approximation of the TAR are also available (Townsend 2003a). This implies that measured amplitude ratios contain integrated astrophysical information about the properties of the stellar atmosphere, notably its aspect-angle dependent effective temperature, gravity, and limb darkening, all of which occur in the theoretical expressions of the ratios. While this was used as a means to try and identify mode degrees from models with fixed input physics prior to space asteroseismology, Gaia now offers the reverse for stars whose mode frequencies and amplitudes are detected in the Gaia BP and RP photometry.
@@
Measured amplitude ratios of stars with already identified modes from mode frequency or period patterns deduced from Kepler, TESS (or future PLATO) light curves thus offer modelling potential. This can be achieved from multi-colour (space) photometry by imposing the mode identification. Since the 23 Dor stars in Fig. 11 have the same solar metallicity and the same mode identification in terms of , the spread in measured amplitude ratios shown in the figure is due to different inclination angles and surface rotation rates, combined with a variety of physical circumstances in their atmosphere. This concerns limb darkening properties, possible latitudinal differential surface rotation, temperature distributions, properties of convection, etc. The treatment of convection has a particularly large impact on the predictions of amplitude ratios and phase differences (Dupret et al. 2005a, b). Thus, exploitation of the amplitude ratios measured with highest precision can in principle allow us to infer which of these atmospheric properties are dominant. This may even help to find optimal stars with a variety of surface properties for calibration purposes of new high-resolution spectrographs capable of detecting time-resolved line-profile variability due to low-amplitude surface velocity fields and/or weak surface spots, including exoplanet hosts (cf. Fig. 1). One can see in Fig. 11 that the uncertainties in the Gaia BP and RP amplitude ratios are currently still too large to exploit this potential, but they should decrease appreciably for the longer light curves to be expected from Gaia Data Releases 4 (2026) and 5. These will occur at the time when the PLATO mission should be operational. PLATO itself offers 2-colour photometry from its two fast cameras and is even more suited to push this research beyond current instrument barriers than Kepler or TESS combined with Gaia. In addition, ground-based survey facilities with capacity to deliver multi-colour photometry at mmag level will become operational soon, such as BlackGEM (Groot et al. 2022) and the Vera Rubin Observatory (Di Criscienzo et al. 2023). The basics of the required modelling tools to interpret the amplitude ratios in moderate and fast rotators were already developed long before the era of space asteroseismology (e.g., Daszyńska-Daszkiewicz et al. 2002; Townsend 2003a). Now, about two decades later, we enter a time where they can be repurposed and applied massively from combined Gaia, Kepler, TESS, PLATO, BlackGEM, and Rubin multicolour survey photometry for all the pulsators with proper identification of their dominant mode(s). Exploitation of the amplitude ratios in terms of internal and atmosphere physics from modern multi-colour photometry should be within reach for modes with amplitudes of several mmag.
8.2 The major need for better theory and modelling tools
@@
Optimal exploitation of the wealth of available and future observations of stellar oscillations requires concerted efforts to develop better theory of non-radial oscillations for rotating single stars and for close binaries, as well as appropriate new modelling tools. In order to treat moderate and fast rotators up to the level of the precision of modern data, we must go beyond the current modelling tools (Vanlaer et al. 2023) and/or use non-linear pulsation theory. Ample detections of amplitude modulation (e.g. Bowman 2016) and non-linear mode coupling (e.g. Huat et al. 2009; Breger et al. 2012; Kurtz et al. 2015; Baade et al. 2016; Van Beeck et al. 2021) are available, while most of current asteroseismic modelling only makes use of measured oscillation frequencies assuming linear modes. Non-linear asteroseismology allows us to exploit mode amplitudes from coupled mode equations instead of relying only on the frequencies of identified modes and assuming that they are linear. Following initial pioneering steps in non-linear space asteroseismology for slowly rotating evolved subdwarfs and white dwarfs by Zong et al. (2016b, a), similar applications to moderate and fast rotators along the main sequence have yet to be tried out. This can be done by extending the principles laid out by Buchler et al. (1997); Goupil et al. (1998); Mourabit & Weinberg (2023) by means of a proper treatment of the rotation. Lee (2012, 2022) and Van Beeck et al. (2023) include the Coriolis force in the sub-inertial regime (cf. Fig. 3), providing enhanced theory to model the numerous gravito-inertial non-linear pulsators with identified three-mode couplings. Asteroseismic modelling based on these non-linear formalisms would allow us to scrutinise transport processes from early phases of evolution onwards. This would constitute a critical step ahead because the effects of angular momentum transport and chemical mixing are cumulative and hence have major implications for the entire life of the star, from birth to death.
@@
Finally, moderate and fast rotators require attention at an even more fundamental level than developing and applying non-linear asteroseismology. Indeed, if we really want to maximise exploitation of the asteroseismic data of such pulsators we must abandon the premise that stars can be approximated as spheres, both at the level of the stellar structure equations and in solving the oscillation equations to compute the modes. Rotating stars are flattened spheroids, requiring at least 2D (for rotating single stars) and optimally 3D (for magnetic spheroids and/or binary systems) treatments, at the level of realistic equilibrium models including the chemical evolution of the star (Deupree 1990, 1995; Rieutord et al. 2016; Mombarg et al. 2023) and of numerical computations of their oscillation modes (Ouazzani et al. 2015; Reese et al. 2021). Major long-term mathematical and computational efforts to achieve this are on the horizon111http://www.4D-STAR.org and offer a fantastic opportunity to bridge the fields of asteroseismology, close binary evolution, binary population synthesis, and gravitational wave astrophysics, as highlighted in the fountain in Fig. 1.
@@
Take away messages from asteroseismology of fast rotators in the literature include: • Asteroseismology of stars rotating up to 70% of their critical Keplerian rotation rate is a mature field of research, delivering their mass, radius, core mass, and core radius with precisions of a few percent for the best cases. • About 80% of the detected g-mode oscillations in intermediate-mass dwarfs are prograde dipole gravito-inertial modes. Such modes deliver the near-core rotation rate to better than 5% accuracy. Values range from zero to Hz. • For a small fraction of the fastest rotators in the mass range M, inertial modes in the convective core couple to gravito-inertial modes in the envelope, offering a direct measurement of the core rotation. • Single intermediate-mass dwarfs rotate almost rigidly, with their core and envelope rotation rate differing by less than 10%. The asteroseismic results imply rapid transport of angular momentum from the core to the surface during the main sequence. • About 15% of the fast rotating dwarfs in the mass range M, as well as some of the outbursting Be pulsators, reveal global Rossby modes. • Nonlinear resonant mode coupling is omnipresent in rotating intermediate-mass pulsators and in their subdwarf and white dwarf successors. This offers a golden opportunity to exploit mode amplitudes in addition to frequencies. • The fastest rotating B and Be pulsators in the mass range M show outbursts in their photometric light curves, often accompanied by non-linear mode coupling. • Asteroseismic core-to-envelope rotation ratios have been determined for a handful of high-mass dwarfs, revealing values between one and five. • The cumulative effect of the measured near-core mixing in dwarfs with a convective core extends the main sequence duration and induces large spreads at the turnoff. This mixing impacts all phases of evolution, notably age-dating, because up to twice as much helium can be produced by the time of the turnoff compared to the case without any near-core mixing. • Evidence for enhanced internal mixing during the core helium burning phase is uncovered from asteroseismology of white dwarfs and subdwarfs, but has yet to be calibrated with high precision. • Measured vertical diffusive envelope mixing levels in intermediate- and high-mass dwarfs range from 10 cm s to 10 cm s. The near-core and envelope mixing profiles are diverse and change over time. • Internal magnetic fields are detected in a fraction of pulsators across the entire evolution and mass ranges accessible by asteroseismology, revealing field strengths between G and about G. • Model-independent dynamical masses of detached eclipsing binaries offer tight constraints on asteroseismic modelling, confirming that stellar models require considerable near-core mixing to bring (core) masses from theory and observations into agreement. • Tidally excited, tidally perturbed, and tidally tilted oscillations occur across a wide range of binaries, including supernova progenitors and members of young open clusters. This makes asteroseismic calibration of tidal dissipation processes and of close binary evolution models possible, including binaries en route to gravitational wave emission.
@@ Acknowledgements.
The authors thank Gaël Buldgen and Jan Henneco for sharing their unpublished work that led to Figures 5 and 9; they are also grateful to Clio Gielen, Siemen Burssens, Jordan Van Beeck, Jan Henneco, and Joris De Ridder for providing Figures 1 and 3, 6, 7, 9, and 10, respectively. Jordan Van Beeck and Vincent Vanlaer are acknowledged for their comments on a draft version of the text prior to submission. We gratefully acknowledge the positive feedback and detailed comments received from two anonymous referees. We also express our appreciation to those colleagues from the community having done the effort to answer our call for comments by offering feedback and suggestions, which helped us to improve the text. The authors acknowledge funding from the KU Leuven Research Council (grant C16/18/005: PARADISE) and from the BELgian federal Science Policy Office (BELSPO) through PRODEX grant PLATO. CA additionally received financial support from the Research Foundation Flanders (FWO) under grant K802922N (Sabbatical leave) and from the European Research Council (ERC) under the Horizon Europe programme (Synergy Grant agreement N101071505: 4D-STAR). Work for this review was partially funded by the European Union. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. The content of this review was fed by inspiring discussions held at the Munich Institute for Astro-, Particle and BioPhysics (MIAPbP), which is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC-2094 – 390783311. CA thanks the organisers and local staff for the invitation and for the good organisation. She is also grateful for the kind hospitality offered by the Center for Computational Astrophysics of the Flatiron Institute of the Simons Foundation in New York, by the Max Planck Institute of Astronomy in Heidelberg, and by the Centre d’Étude Atomaire in Saclay during her sabbatical work visits in academic year 2022 – 2023.References
@@
- Bellinger et al. (2024) Bellinger, E. P., de Mink, S. E., van Rossem, W. E., & Justham, S. 2024, ApJ, submitted, arXiv:2311.00038
@@
- Buldgen et al. (2023a) Buldgen, G., Eggenberger, P., Noels, A., et al. 2023a, A&A, 669, L9
- Buldgen et al. (2023b) Buldgen, G., Noels, A., Baturin, V. A., et al. 2023b, A&A, in press, arXiv:2308.13368
@@
- Domiciano de Souza et al. (2014a) Domiciano de Souza, A., Kervella, P., Moser Faes, D., et al. 2014a, A&A, 569, A10
- Domiciano de Souza et al. (2014b) Domiciano de Souza, A., Kervella, P., Moser Faes, D., et al. 2014b, A&A, 569, A10
@@
- Felce & Fuller (2023) Felce, C. & Fuller, J. 2023, MNRAS, in press [arXiv:2309.08506]
- Fritzewski et al. (2024) Fritzewski, D. J., Aerts, C., Mombarg, J. S. G., Gossage, S., & Van Reeth, T. 2024, A&A, in press, arXiv:2402.05168
- Fritzewski et al. (2023) Fritzewski, D. J., Van Reeth, T., Aerts, C., et al. 2023, A&A, in press, arXiv:2310.18426
@@
- Groot et al. (2022) Groot, P. J., Bloemen, S., Vreeswijk, P. M., et al. 2022, in Society of Photo-Optical Instrumentation Engineers (SPIE) Conference Series, Vol. 12182, Ground-based and Airborne Telescopes IX, ed. H. K. Marshall, J. Spyromilio, & T. Usuda, 121821V
@@
- Jennings et al. (2023) Jennings, Z., Southworth, J., Pavlovski, K., & Van Reeth, T. 2023, A&A, in press, arXiv:2311.02064
@@
- Li et al. (2023a) Li, G., Deheuvels, S., Li, T., Ballot, J., & Lignières, F. 2023a, A&A, in press, arXiv:2309.13756
@@
- Lundkvist et al. (2024) Lundkvist, M. S., Kjeldsen, H., Bedding, T. R., McCaughrean, M. J., & many others. 2024, ApJ, in press
@@
- Marchant & Bodensteiner (2024) Marchant, P. & Bodensteiner, J. 2024, ARA&A, submitted, arXiv:2311.01865
@@
- Rauer et al. (2014) Rauer, H., Catala, C., Aerts, C., et al. 2014, Experimental Astronomy, 38, 249
@@
- Rui et al. (2023) Rui, N. Z., Ong, J. M. J., & Mathis, S. 2023, MNRAS, submitted, arXiv:2310.19873
@@
- Van Beeck et al. (2023) Van Beeck, J., Van Hoolst, T., Aerts, C., & Fuller, J. 2023, A&A, submitted, arXiv:2311.02972

### v4

@@ 8.1 Pushing Gaia beyond its limits to assist in the asteroseismology of intermediate- and high-mass stars
Pushing even more beyond the limits, we can now fully integrate the knowledge from Gaia’s three-band and Kepler’s and/or TESS’s (or in the future PLATO’s) single-band light curves to improve asteroseismic modelling. Such a synergistic integration of data is illustrated in Fig. 11, which compares the amplitudes of the dominant dipole prograde mode in the Gaia BP and Kepler bands for 309 Dor stars. Theory predicts the amplitudes of dipole modes in Dor pulsators to be higher at the BP than at the wavelengths for Dor pulsators (Aerts et al. 2004). This is in agreement with the results for g-mode pulsators shown in the left panel in Fig. 11. This panel shows the amplitude ratios of the 23 Dor stars among the sample of 309, for which a harmonic fit with the dominant Kepler frequency explains at least 50% of the overall variance in the Gaia BP and RP time series. Measured amplitudes depend on the inclination angle of the star’s symmetry axis of the oscillations with respect to the line of sight and on the position of the surface nodal lines. However, for slow rotators, both dependencies drop out of the theoretical expression of amplitude ratios with respect to a reference wavelength (see Chapter 6, Aerts et al. 2010, for details). Theoretical expressions for the ratios in the approximation of the TAR are also available (Townsend 2003a). This implies that measured amplitude ratios contain integrated astrophysical information about the properties of the stellar atmosphere, notably its aspect-angle dependent effective temperature, gravity, and limb darkening, all of which occur in the theoretical expressions of the ratios. While this was used as an attempt to identify mode degrees from models with fixed input physics prior to space asteroseismology, Gaia now also offers the reverse for stars whose frequencies and amplitudes of already identified modes are detected in the Gaia BP and RP photometry.
In addition, Fritzewski et al. (2024c) showed that the Gaia BP and RP photometry may lead to the identification of the degree of low-order modes in Cep stars. They designed a novel probabilistic identification method and applied it to the dominant mode detected consistently in the Gaia BP, RP, and TESS data of 164 Cep stars included in Fig. 10, among which 121 new pulsators discovered from Gaia. This mode identification and the detection of rotationally split modes in the TESS FFI data led to the first forward modelling of a population of high-mass pulsators, resulting in masses between 9 M⊙ and about 20 M⊙ for these 164 stars. For 48 of them, limits on the core-to-surface rotation ratios were derived from the combined Gaia and TESS data; most of these stars have values below 2, but a few objects reveal a ratio up to 4. Such internal rotation estimates were only available for five class members before Gaia DR3.
@@
These examples illustrate that Gaia already has an important role to play in asteroseismic modelling of intermediate- and high-mass pulsators. Much more is to come in the near future. Since the 23 Dor stars in Fig. 11 have similar metallicity and the same mode identification in terms of , future exploitation of their amplitude ratios may reveal their atmospheric properties, which are currently unknown. This may even help us to find optimal stars with a variety of surface phenomena, such as spots and lattitudinal-differential rotation, for calibration purposes of new high-resolution spectrographs to study exoplanet hosts (see Fig. 1). Figure 11 shows that the uncertainties in the Gaia BP and RP amplitude ratios are currently still too large to exploit this potential, but they should decrease appreciably for the longer light curves to be expected from Gaia Data Releases 4 (2026) and 5. These will occur at the time when the PLATO mission should be operational. PLATO itself offers two-colour photometry from its two fast cameras, and is even more suited to push this research beyond current instrument barriers than Kepler or TESS combined with Gaia. In addition, ground-based survey facilities with the capacity to deliver multi-colour photometry at the millimagnitude level are operational or will become operational soon, such as BlackGEM (Groot et al. 2024) and the Vera Rubin Observatory (formerly known as the Large Synoptic Survey Telescope (LSST), Di Criscienzo et al. 2023). The basics of modelling tools used to interpret the amplitude ratios in moderate and fast rotators had already been developed long before the era of space asteroseismology (e.g. Daszyńska-Daszkiewicz et al. 2002; Townsend 2003a). Now, about two decades later, modern tools such as those designed by Fritzewski et al. (2024c) can be applied massively from combined Gaia, Kepler, TESS, PLATO, BlackGEM, and Rubin multi-colour survey photometry for all the pulsators with proper identification of their mode(s). Exploitation of the amplitude ratios in terms of internal and atmosphere physics from modern multi-colour photometry should be within reach for modes with amplitudes of several mmag.
8.2 The need for better theory and modelling tools
@@
Optimal exploitation of the wealth of available and future observations of stellar oscillations requires concerted efforts to develop better theory. This need is particularly prominent for the theory and modelling of non-radial oscillations in fast-rotating single stars and close binaries. In order to treat moderate and fast rotators up to the level of the precision of modern data, we must go beyond the current modelling tools and develop non-linear pulsation theory involving mode coupling and non-linear inversion tools (Vanlaer et al. 2023; Farrell et al. 2024). Ample detections of amplitude modulation (e.g. Bowman 2016) and non-linear mode coupling (e.g. Huat et al. 2009; Breger et al. 2012; Kurtz et al. 2015; Baade et al. 2016; Van Beeck et al. 2021) in fast rotators are available, while most asteroseismic modelling currently only relies on measured frequencies assuming linear and uncoupled modes. Non-linear asteroseismology would allow us to exploit mode amplitudes from coupled mode equations. Following the initial pioneering steps in non-linear space asteroseismology for slow-rotating evolved sub-dwarfs and white dwarfs by Zong et al. (2016b, a), similar applications to moderate and fast rotators along the main sequence have yet to be tried out. This can be done by extending the principles laid out by Buchler et al. (1997); Goupil et al. (1998); Mourabit & Weinberg (2023), among others, by means of a proper treatment of the rotation. Lee (2012, 2022) and Van Beeck et al. (2024) include the Coriolis force in the sub-inertial regime (see Fig. 3), providing enhanced theory to model the numerous gravito-inertial non-linear pulsators with identified multi-mode couplings. Asteroseismic modelling based on these non-linear formalisms would allow us to scrutinise transport processes from early phases of evolution onwards. This would constitute a critical step ahead because the effects of angular momentum transport and chemical mixing are cumulative, and hence have major implications for the entire life of the star, from birth to death.
@@
Finally, moderate and fast rotators require attention at an even more fundamental level than developing and applying non-linear asteroseismology. If we truly want to maximise the exploitation of the asteroseismic data of such pulsators, we must abandon the premise that stars can be approximated as spheres, both at the level of the stellar structure equations and in solving the oscillation equations to compute the modes. Rotating stars are oblate spheroids, requiring at least 2D (for rotating single stars) and optimally 3D (for magnetic spheroids and/or binary systems) treatments. Realistic evolution models including the chemical evolution of stars are being constructed (Deupree 1990, 1995; Rieutord et al. 2016; Mombarg et al. 2023) and so are numerical computations of their oscillation modes (Ouazzani et al. 2015; Reese et al. 2021). Major long-term mathematical and computational efforts to achieve this in a systematic approach are on the horizon111www.4D-STAR.org and offer a fantastic opportunity to bridge the fields of asteroseismology, close binary evolution, binary population synthesis, and gravitational wave astrophysics, as highlighted in the fountain in Fig. 1.
@@
Takeaway messages from the asteroseismology of fast rotators in the literature include: • Asteroseismology of stars rotating up to 70% of their critical Keplerian rotation rate is a mature field of research, delivering their mass, radius, core mass, and core radius with a precision of a few percent for the best cases. • About 80% of the detected g-mode oscillations in intermediate-mass dwarfs are prograde dipole gravito-inertial modes. These modes deliver the near-core rotation frequency to better than 5% accuracy. Values range from zero to Hz. • For a small fraction of the fastest rotators in the mass range M⊙, inertial modes in the convective core couple to gravito-inertial modes in the envelope, offering a direct measurement of the rotating core properties. • Single intermediate-mass dwarfs have modest radial-differential rotation, with their core and surface rotation frequencies differing by less than 10%. The asteroseismic results imply rapid transport of angular momentum from the core to the surface during the main sequence. • About 15% of the fast-rotating dwarfs in the mass range M⊙, as well as some of the outbursting Be pulsators, reveal global Rossby modes. • Non-linear resonant mode coupling is omnipresent in rotating intermediate-mass pulsators and in their sub-dwarf and white dwarf successors. This offers a golden opportunity to exploit mode amplitudes in addition to frequencies. • The fastest rotating B and Be pulsators in the mass range M⊙ show outbursts in their photometric light curves, often accompanied by non-linear mode coupling. • Asteroseismic core-to-surface rotation ratios have been determined for about 50 high-mass dwarfs in the mass range M⊙, revealing values below 2 for most of them, but a value up to 5 for a few stars. • The cumulative effect of the measured near-core mixing in dwarfs with a convective core extends the duration of their main sequence and induces large spreads at the turnoff. This mixing impacts all phases of evolution, notably age-dating and chemical yield production. Up to twice as much helium can be produced by the time of the turnoff compared to the case without any near-core mixing. • Evidence of enhanced internal mixing during the core-helium burning phase is uncovered from asteroseismology of white dwarfs and sub-dwarfs, but has yet to be calibrated with high precision for large samples. • Measured vertical diffusive envelope mixing levels in single intermediate- and high-mass dwarfs range from 10 cm2 s-1 to 107 cm2 s-1. The near-core and envelope mixing profiles are diverse and change over time. • Internal magnetic fields are detected in a fraction of pulsators across the entire evolution and mass ranges accessible by asteroseismology, revealing field strengths between G and about G. • Model-independent dynamical masses of detached eclipsing binaries offer tight constraints on asteroseismic modelling, confirming that stellar models require near-core boundary mixing to bring (core) masses from theory and observations into agreement. • Tidally excited, tidally perturbed, and tidally tilted oscillations occur across a wide range of binaries, including supernova progenitors and members of young open clusters. This makes asteroseismic calibration of tidal dissipation processes and of close binary evolution models possible, including binaries en route to gravitational wave emission.
@@ Acknowledgements.
The authors thank Gaël Buldgen and Jan Henneco for sharing their work that led to Figures 5 and 9; they are also grateful to Clio Gielen, Siemen Burssens, Jordan Van Beeck, Jan Henneco, and Joris De Ridder for providing Figures 1 and 3, 6, 7, 9, and 10, respectively. Jordan Van Beeck and Vincent Vanlaer are acknowledged for their comments on a draft version of the text prior to submission. We gratefully acknowledge the positive feedback and detailed comments received from two anonymous referees. We also express our appreciation to those colleagues from the community having done the effort to answer our call for comments by offering feedback and suggestions, which helped us to improve the text. The authors acknowledge funding from the KU Leuven Research Council (grant C16/18/005: PARADISE) and from the BELgian federal Science Policy Office (BELSPO) through PRODEX grant PLATO. CA additionally received financial support from the Research Foundation Flanders (FWO) under grant K802922N (Sabbatical leave) and from the European Research Council (ERC) under the Horizon Europe programme (Synergy Grant agreement N∘101071505: 4D-STAR). Work for this review was partially funded by the European Union. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. The content of this review was fed by inspiring discussions held at the Munich Institute for Astro-, Particle and BioPhysics (MIAPbP), which is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC-2094 – 390783311. CA thanks the organisers and local staff for the invitation and for the good organisation. She is also grateful for the kind hospitality offered by the Center for Computational Astrophysics of the Flatiron Institute of the Simons Foundation in New York, by the Max Planck Institute of Astronomy in Heidelberg, and by the Centre d’Étude Atomaire in Saclay during her sabbatical work visits in academic year 2022 – 2023.References
@@
- Bellinger et al. (2024) Bellinger, E. P., de Mink, S. E., van Rossem, W. E., & Justham, S. 2024, ApJ, 967, L39
@@
- Brinkman et al. (2024) Brinkman, H. E., Roberti, L., Kemp, A., et al. 2024, A&A, in press, arXiv:2406.02404
@@
- Buldgen et al. (2023) Buldgen, G., Eggenberger, P., Noels, A., et al. 2023, A&A, 669, L9
- Buldgen et al. (2024) Buldgen, G., Noels, A., Baturin, V. A., et al. 2024, A&A, 681, A57
@@
- de Laverny et al. (2024) de Laverny, p., Recio-Blanco, A., Aerts, C., & Palicio, P. A. 2024, A&A, under review
@@
- Domiciano de Souza et al. (2014) Domiciano de Souza, A., Kervella, P., Moser Faes, D., et al. 2014, A&A, 569, A10
@@
- Farrell et al. (2024) Farrell, E., Buldgen, G., Meynet, G., et al. 2024, A&A, 686, A267
- Felce & Fuller (2023) Felce, C. & Fuller, J. 2023, MNRAS, 526, 6168
- Fritzewski et al. (2024a) Fritzewski, D. J., Aerts, C., Mombarg, J. S. G., Gossage, S., & Van Reeth, T. 2024a, A&A, 684, A112
- Fritzewski et al. (2024b) Fritzewski, D. J., Van Reeth, T., Aerts, C., et al. 2024b, A&A, 681, A13
- Fritzewski et al. (2024c) Fritzewski, D. J., Vanrespaille, M., Aerts, C., Hey, D., & De Ridder, J. 2024c, A&A, submitted, arXiv:2408.06097
@@
- Groot et al. (2024) Groot, P. J., Bloemen, S., Vreeswijk, P., et al. 2024, PASP, submitted, arXiv:2405.18923
@@
- Guo et al. (2024) Guo, Z., Bedding, T. R., Pamyatnykh, A. A., et al. 2024, MNRAS, under review, arXiv:2406.15678
@@
- Henneco et al. (2024) Henneco, J., Schneider, F. R. N., Hekker, S., & Aerts, C. 2024, A&A, in press, arXiv:2406.14416
@@
- Hirschi et al. (2004) Hirschi, R., Meynet, G., & Maeder, A. 2004, A&A, 425, 649
- Hirschi et al. (2005) Hirschi, R., Meynet, G., & Maeder, A. 2005, A&A, 433, 1013
@@
- Jennings et al. (2024) Jennings, Z., Southworth, J., Pavlovski, K., & Van Reeth, T. 2024, MNRAS, 527, 4052
@@
- Kemp et al. (2024) Kemp, A., Tkachenko, A., Torres, G., et al. 2024, A&A, in press, arXiv:2406.04131
@@
- Li et al. (2023a) Li, G., Deheuvels, S., Li, T., Ballot, J., & Lignières, F. 2023a, A&A, 680, A26
@@
- Lundkvist et al. (2024) Lundkvist, M. S., Kjeldsen, H., Bedding, T. R., et al. 2024, ApJ, 964, 110
@@
- Marchant & Bodensteiner (2024) Marchant, P. & Bodensteiner, J. 2024, ARA&A, in press, arXiv:2311.01865
@@
- Moe & Di Stefano (2017) Moe, M. & Di Stefano, R. 2017, ApJS, 230, 15
- Mombarg et al. (2024a) Mombarg, J. S. G., Aerts, C., & Molenberghs, G. 2024a, A&A, 685, A21
- Mombarg et al. (2024b) Mombarg, J. S. G., Aerts, C., Van Reeth, T., & Hey, D. 2024b, A&A, under review
@@
- Rauer et al. (2024) Rauer, H., Aerts, C., Cabrera, J., et al. 2024, Experimental Astronomy, under review, arXiv:2406.05447
@@
- Rui & Fuller (2024) Rui, N. Z. & Fuller, J. 2024, Open Journal of Astrophysics, submitted, arXiv:2404.14474
- Rui et al. (2024) Rui, N. Z., Ong, J. M. J., & Mathis, S. 2024, MNRAS, 527, 6346
@@
- Van Beeck et al. (2024) Van Beeck, J., Van Hoolst, T., Aerts, C., & Fuller, J. 2024, A&A, 687, A265